# The twisted coarse Baum–Connes conjecture and relative hyperbolic groups

### v1

@@ 5.1 Decomposition
Therefore, the family of triples $(B^{\omega};I^{\omega},J^{\omega})_{\omega}$ is uniformly excisive. ∎
Moreover, for $s\geq s_{0}$, define
$B_{r}^{\omega,s}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0}^{+r});\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}+C^{*}_{L,0}(P_{s_{0}}(Y_{1}^{+r});\Gamma(Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}$
$\quad\quad\quad\quad+C^{*}_{L,0}(P_{s_{0}}(Y_{0}^{+r})\cap P_{s_{0}}(Y_{1}^{+r}));\Gamma(Y_{0}^{+r}\cap Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{sr}$
$I_{r}^{\omega,s}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0}^{+r});\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}+C^{*}_{L,0}(P_{s}(Y_{0}^{+r}\cap Y_{1}^{+r});\Gamma(Y_{0}^{+r}\cap Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{sr}$
$J_{r}^{\omega,s}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{1}^{+r});\Gamma(Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}+C^{*}_{L,0}(P_{s}(Y_{0}^{+r}\cap Y_{1}^{+r});\Gamma(Y_{0}^{+r}\cap Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{sr}$
define a $C^{*}$-subalgebra $B^{\omega,s}:=\overline{\bigcup_{r\geq 0}B_{r}^{\omega,s}}$ of $C^{*}_{L,0}(P_{s_{0}};\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ and two ideals $I^{\omega,s}:=\overline{\bigcup_{r\geq 0}I_{r}^{\omega,s}}$ and $J^{\omega,s}:=\overline{\bigcup_{r\geq 0}J_{r}^{\omega,s}}$ of $B^{\omega,s}$.
@@
The family of triples $(I^{\omega,s}(X),J^{\omega,s}(X);B^{\omega,s}(X))_{\omega,s}$ is uniformly excisive, where $\omega$ runs all decompositions $\omega=(Y_{1},Y_{2})$ of $X$ and $s$ runs all $s\geq s_{0}$.
@@
Same as the proof of Lemma 5.1. ∎
We have a preliminary lemma (analogous to [GWY24, Lemma 7.14]).
Lemma 5.3.
Let $X$ be a metric space and $r_{0},s_{0}>0$. Assume that $X$ admits a decomposition $X=Y_{0}\cup Y_{1}$. Then we have
$C^{*}_{L,0}(P_{s_{0}}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))_{r_{0}}\subset C^{*}_{L,0}$ $\left(P_{s_{0}}(Y_{0}^{+(2s_{0}+r_{0})});\Gamma(Y_{0}^{+(2s_{0}+r_{0})},\mathcal{A}\otimes\mathcal{K})\right)_{r_{0}}$
$\quad\quad+C^{*}_{L,0}\left(P_{s_{0}}(Y_{1}^{+(2s_{0}+r_{0})}),\Gamma(Y_{1}^{+(2s_{0}+r_{0})};\mathcal{A}\otimes\mathcal{K})\right)_{r_{0}}.$
Proof.
We have
$P_{s_{0}}(X)\subset P_{s_{0}}\left(Y_{0}^{+s_{0}}\right)\cup P_{s_{0}}\left(Y_{1}^{+s_{0}}\right).$
We denote by $\chi_{0}$ the characteristic function of $N_{r_{0}}(P_{s_{0}}(Y_{0}^{+s_{0}}))\subset P_{s_{0}}(X)$. For any $A\in C^{*}_{L,0}(P_{s_{0}}(X))_{r_{0}}$, we define
$A_{0}=\chi_{0}A\chi_{0},~A_{1}=A-A_{0}.$
Then $A_{0}$ is supported on $N_{r_{0}}(P_{s_{0}}(Y_{0}^{+s_{0}}))\times N_{r_{0}}(P_{s_{0}}(Y_{0}^{+s_{0}}))$ and
$A_{1}=A(1-\chi_{0})+(1-\chi_{0})A\chi_{0}$
is supported on $N_{r_{0}}(P_{s_{0}}(Y_{1}^{+s_{0}}))\times N_{r_{0}}(P_{s_{0}}(Y_{1}^{+s_{0}}))$ since the propagation of $A$ is less than $r_{0}$. So we have
$C^{*}_{L,0}(P_{s_{0}}(X))_{r_{0}}\subset C^{*}_{L,0}\left(N_{r_{0}}\left(P_{s_{0}}\left(Y_{0}^{+s_{0}}\right)\right)\right)_{r_{0}}+C^{*}_{L,0}\left(N_{r_{0}}\left(P_{s_{0}}\left(Y_{1}^{+s_{0}}\right)\right)\right)_{r_{0}}.$
But for $i=0,1$, we have
$N_{r_{0}}\left(P_{s_{0}}\left(Y_{i}^{+s_{0}}\right)\right)\subset P_{s_{0}}\left(Y_{i}^{+(2s_{0}+r_{0})}\right).$
Therefore we have
$C^{*}_{L,0}(P_{s_{0}}(X))_{r_{0}}\subset C^{*}_{L,0}\left(P_{s_{0}}\left(Y_{0}^{+(2s_{0}+r_{0})}\right)\right)_{r_{0}}+C^{*}_{L,0}\left(P_{s_{0}}\left(Y_{1}^{+(2s_{0}+r_{0})}\right)\right)_{r_{0}}.$
As well as the first half of the proof of Lemma 5.1, if $A\in C^{*}_{L,0}(P_{s_{0}}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))_{r_{0}}$, we have
$A_{i}\in C^{*}_{L,0}\left(P_{s_{0}}(Y_{i}^{+(2s_{0}+r_{0})});\Gamma(Y_{0}^{+(2s_{0}+r_{0})},\mathcal{A}\otimes\mathcal{K})\right)_{r_{0}}$
for $i=0,1$ again since they are restrictions. ∎
@@ 5.2 Reduction to $(B(n),d_{S})$
In this section, we prove Lemma 4.8. Let $x\in K_{*}^{r_{0},\frac{1}{8}}(\mathbb{C}_{L,0}(P_{s_{0}}(\bigsqcup X);\Gamma(\bigsqcup X,\mathcal{A}\otimes\mathcal{K}))$ be any element for any $r_{0}$ and $s_{0}$, where $\Gamma\left(\bigsqcup X,\mathcal{A}\otimes\mathcal{K}\right)$ is any stable coarse $\left(\bigsqcup_{X\in\mathcal{X}}X\right)$-algebra. It suffices to show that there exists $s\geq s_{0}$ such that $x=0$ in
@@
For any subset $X^{\prime}\subset\bigsqcup_{X\in\mathcal{X}}X$, $\Gamma(X^{\prime},\mathcal{A}\otimes\mathcal{K})$ is the coarse $X^{\prime}$-algebra defined in (9). Decompose
@@
such that the family $\{Y_{i}^{(j)}(X)\}_{j}$ is $3(2r_{0}+s_{0})$-separated and each $Y_{i}^{(j)}(X)$ is in $\mathcal{Y}$ for $i=0,1$ for some $\mathcal{Y}\in\mathfrak{D}$. Then this gives a decomposition
$\bigsqcup_{X\in\mathcal{X}}X=\bigsqcup_{j,X}Y_{0}^{(j)}(X)\cup\bigsqcup_{j,X}Y_{1}^{(j)}(X).$
@@
We denote this decomposition of $\bigsqcup_{X\in\mathcal{X}}X$ by $\omega=\left(\bigsqcup_{j,X}Y_{0}^{(j)}(X),\bigsqcup_{j,X}Y_{1}^{(j)}(X)\right)$ and denote $Y_{i}:=\bigsqcup_{j,X}Y_{i}^{(j)}(X)$ where the distance between distinct components is taken to be infinite. By Lemma 5.3, $C^{*}_{L,0}\left(P_{s_{0}}(\bigsqcup X);\Gamma(\bigsqcup X,\mathcal{A}\otimes\mathcal{K})\right)_{r_{0}}$ is contained in
@@
$C^{*}_{L,0}\left(P_{s_{0}}\left(Y_{0}^{+(2s_{0}+r_{0})}\right);\Gamma\left(Y_{0}^{+(2s_{0}+r_{0})},\mathcal{A}\otimes\mathcal{K}\right)\right)_{r_{0}}+C^{*}_{L,0}\left(P_{s_{0}}\left(Y_{0}^{+(2s_{0}+r_{0})}\right);\Gamma\left(Y_{0}^{+(2s_{0}+r_{0})},\mathcal{A}\otimes\mathcal{K}\right)\right)_{r_{0}},$
@@
which is contained in $B_{2s_{0}+r_{0}}^{\omega}$. Therefore, $x$ is sent to an element in $K^{\frac{1}{8},2s_{0}+r_{0}}(B^{\omega})$. There exists $r_{1}\geq 2s_{0}+r_{0}$ and the boundary map $\partial$, which fits in the commutative diagram below for every $s_{1}\geq s_{0}$:
$K_{*}^{2r_{0}+s_{0},\frac{1}{8}}(B^{\omega})$$K_{*}^{r_{1},\frac{1}{8}}(I^{\omega}\cap J^{\omega})$$K_{*}^{2r_{0}+s_{0},\frac{1}{8}}(B^{\omega,s_{1}})$$K_{*}^{r_{1},\frac{1}{8}}(I^{\omega,s_{1}}\cap J^{\omega,s_{1}}).$$\scriptstyle{\partial}\scriptstyle{\partial}$
Since $\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K})\cap\Gamma(Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K})=\Gamma(Y_{0}^{+r}\cap Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K})$, the right vertical map can be identified with
$K_{*}^{r_{1},\frac{1}{8}}(C^{*}_{L,0}$ $(P_{s_{0}}(Y_{0}^{+r_{1}}\cap Y_{1}^{+r_{1}});\Gamma(Y_{0}^{+r_{1}}\cap Y_{1}^{+r_{1}},\mathcal{A}\otimes\mathcal{K}))$
@@
$\longrightarrow K_{*}^{r_{1}s_{1},\frac{1}{8}}(C^{*}_{L,0}(P_{s_{1}}(Y_{0}^{+r_{1}}\cap Y_{1}^{+r_{1}});\Gamma(Y_{0}^{+r_{1}}\cap Y_{1}^{+r_{1}},\mathcal{A}\otimes\mathcal{K}))).$
@@
By the coarse invariance and the permanence under taking subspace of the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras, $Y_{0}^{+r_{1}}\cap Y_{1}^{+r_{1}}$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras. Therefore, $\partial x$ is sent to 0 for sufficiently large $s_{1}$. By the controlled-exactness of the Mayer-Vietoris sequence, there exists $r_{2}$ and
$y\in K_{*}^{r_{2},\frac{1}{8}}(I^{\omega,s_{1}}),~z\in K_{*}^{r_{2},\frac{1}{8}}(J^{\omega,s_{1}})$
such that $x=y+z$ in $K_{*}^{r_{2},\frac{1}{8}}(B^{\omega,s_{1}})$. Since
$I^{\omega,s_{1}}_{r_{2}}\subset I^{\omega}_{s_{1}r_{2}},$
by the inductive assumption, $y=0$ in $K_{*}^{s,\frac{1}{8}}(I^{\omega})$ for sufficiently large $s$ and the same thing holds for $z$. Therefore, we have

### v2

@@ 5.1 Decomposition
For any partial translation $v$ on $Y_{i}$, the functions $(A_{0}\mu_{1})^{v}$ and $(A_{1}\mu_{0})^{v}$ is in $\Gamma(Y_{i},\mathcal{A}\otimes\mathcal{K})$ and so we have $B\in I^{X,R,\omega}_{r_{0}+1}\cap J^{X,R,\omega}_{r_{0}+1}$. Therefore, the family of triples $(I^{X,R,\omega},J^{X,R,\omega};B^{X,R,\omega})_{X,R,\omega}$ is uniformly excisive. ∎
Moreover, for $s\geq s_{0}$, we take a net
$Z_{r}:=P_{s}\left(\bigsqcup_{j}Y_{0}^{(j)}\right)^{+r}\cap P_{s}\left(\bigsqcup_{j}Y_{1}^{(j)}\right)^{+r}\cap X$
of $P_{s}\left(\bigsqcup_{j}Y_{0}^{(j)}\right)^{+r}\cap P_{s}\left(\bigsqcup_{j}Y_{1}^{(j)}\right)^{+r}$. Then we define
$B_{r}^{X,R,\omega,s}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0})^{+r};\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K}))_{r}+C^{*}_{L,0}(P_{s_{0}}(Y_{1})^{+r};\Gamma(Y_{1},\mathcal{A}\otimes\mathcal{K}))_{r}$
$\quad\quad\quad\quad+C^{*}_{L,0}\left(P_{s}\left(\bigsqcup_{j}Y_{0}^{(j)}\right)^{+r}\cap P_{s}\left(\bigsqcup_{j}Y_{1}^{(j)}\right)^{+r};\Gamma(Z_{r},\mathcal{A}\otimes\mathcal{K})\right)_{sr},$
$I_{r}^{X,R,\omega,s}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0})^{+r};\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K}))_{r}$
$\quad\quad\quad\quad+C^{*}_{L,0}\left(P_{s}\left(\bigsqcup_{j}Y_{0}^{(j)}\right)^{+r}\cap P_{s}\left(\bigsqcup_{j}Y_{1}^{(j)}\right)^{+r};\Gamma(Z_{r},\mathcal{A}\otimes\mathcal{K})\right)_{sr},$
$J_{r}^{X,R,\omega,s}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{1})^{+r};\Gamma(Y_{1},\mathcal{A}\otimes\mathcal{K}))_{r}$
$\quad\quad\quad\quad+C^{*}_{L,0}\left(P_{s}\left(\bigsqcup_{j}Y_{0}^{(j)}\right)^{+r}\cap P_{s}\left(\bigsqcup_{j}Y_{1}^{(j)}\right)^{+r};\Gamma(Z_{r},\mathcal{A}\otimes\mathcal{K})\right)_{sr},$
where the Rips complex $P_{s}\left(\bigsqcup_{j}Y_{i}^{(j)}\right)$ is taken in terms of the separated coarse disjoint union $\bigsqcup_{j}Y_{i}^{(j)}$ and the $r$-neighborhood $P_{k}\left(\bigsqcup_{j}Y_{i}^{(j)}\right)^{+r}$ is taken relative to $P_{s_{0}}(X)$, i.e.
$P_{s}\left(\bigsqcup_{j}Y_{0}^{(j)}\right)^{+r}=P_{s}\left(\bigsqcup_{j}Y_{0}^{(j)}\right)\cup\left\{p\in P_{s_{0}}(X);d\left(p,P_{s}\left(\bigsqcup_{j}Y_{0}^{(j)}\right)\cap P_{s_{0}}(X)\right)\leq r\right\}.$ (10)
Then define a $C^{*}$-algebras $B^{X,R,\omega,s}:=\overline{\bigcup_{r\geq 0}B_{r}^{X,R,\omega,s}}$ and two ideals $I^{X,R,\omega,s}:=\overline{\bigcup_{r\geq 0}I_{r}^{X,R,\omega,s}}$ and $J^{X,R,\omega,s}:=\overline{\bigcup_{r\geq 0}J_{r}^{X,R,\omega,s}}$ of $B^{X,R,\omega,s}$.
@@
The family $\left(I^{X,R,\omega,s},J^{X,R,\omega,s};B^{X,R,\omega,s}\right)_{X,R,\omega,s}$ is uniformly excisive, where $X$ runs over all metric spaces with bounded geometry, $R$ runs over all increasing functions of $r$ satisfying $R_{r}>3(k_{0}+1)r$, $\omega$ runs over all decompositions $\omega=(Y_{0},Y_{1})$ of $X$ and $s$ runs over all numbers $s\geq s_{0}$.
@@
This follows from the same arguments in the proof of Lemma 5.1. ∎
@@ 5.2 Reduction to $(B(n),d_{S})$
In this section, we prove Lemma 4.8. Let $x\in K_{*}^{r_{0},\frac{1}{8}}(C^{*}_{L,0}(P_{s_{0}}(\bigsqcup X);\Gamma(\bigsqcup X,\mathcal{A}\otimes\mathcal{K}))$ be any element for any $r_{0}$ and $s_{0}$, where $\Gamma\left(\bigsqcup X,\mathcal{A}\otimes\mathcal{K}\right)$ is any stable coarse $\left(\bigsqcup_{X\in\mathcal{X}}X\right)$-algebra. It suffices to show that there exists $s\geq s_{0}$ such that $x=0$ in
@@
We denote $\tilde{X}=\bigsqcup_{X\in\mathcal{X}}X$. For any subset $X^{\prime}\subset\tilde{X}$, recall that $\Gamma(X^{\prime},\mathcal{A}\otimes\mathcal{K})$ is the coarse $X^{\prime}$-algebra defined in (9). We regard the $r_{2}$ of Lemma 3.10 for the uniformly excisive pair of Lemma 5.2 as a function of $r_{0}$ and define a function $R:r_{0}\mapsto(3s_{0}+1)r_{2}$. Decompose
@@
such that the family $\{Y_{i}^{(j)}(X)\}_{j}$ is $3R_{r_{0}}$-separated and each $Y_{i}^{(j)}(X)$ is in $\mathcal{Y}$ for $i=0,1$ for some $\mathcal{Y}\in\mathfrak{D}$. Then this gives a decomposition
$\tilde{X}=\bigsqcup_{j,X}Y_{0}^{(j)}(X)\cup\bigsqcup_{j,X}Y_{1}^{(j)}(X).$
@@
We denote this decomposition of $\tilde{X}$ by $\omega=\left(\bigsqcup_{j,X}Y_{0}^{(j)}(X),\bigsqcup_{j,X}Y_{1}^{(j)}(X)\right)$ and denote $Y_{i}:=\bigsqcup_{j,X}Y_{i}^{(j)}(X)$ where the distance between distinct components is taken to be infinite. Note that $C^{*}_{L,0}\left(P_{s_{0}}(\bigsqcup X);\Gamma(\bigsqcup X,\mathcal{A}\otimes\mathcal{K})\right)_{r_{0}}$ is contained in
@@
$C^{*}_{L,0}\left(P_{s_{0}}(Y_{0})^{+r_{0}};\Gamma\left(Y_{0},\mathcal{A}\otimes\mathcal{K}\right)\right)_{r_{0}}+C^{*}_{L,0}\left(P_{s_{0}}(Y_{1})^{+r_{0}};\Gamma\left(Y_{1},\mathcal{A}\otimes\mathcal{K}\right)\right)_{r_{0}}\subset B_{r_{0}}^{\tilde{X},R,\omega}.$
@@
Therefore, $x$ can be regarded as an element in $K^{\frac{1}{8},r_{0}}(B^{\tilde{X},R,\omega})$. There exists $r_{1}\geq r_{0}$ and the boundary maps $\partial$, which fits in the commutative diagram below for every $s_{1}\geq s_{0}$:
$K_{*}^{r_{0},\frac{1}{8}}(B^{\tilde{X},R,\omega})$$K_{*}^{r_{1},\frac{1}{8}}(I^{\tilde{X},R,\omega}\cap J^{\tilde{X},R,\omega})$$K_{*}^{r_{0},\frac{1}{8}}(B^{\tilde{X},R,\omega,s_{1}})$$K_{*}^{(s_{0}+1)r_{1},\frac{1}{8}}(I^{\tilde{X},R,\omega,s_{1}}\cap J^{\tilde{X},R,\omega,s_{1}}).$$\scriptstyle{\partial}\scriptstyle{\partial}$
Note that since we have
$P_{s_{0}}(N_{r}(Y_{i}(X)))\subset P_{s_{0}}(Y_{i}(X))^{+r}\subset P_{s_{0}}(N_{(s_{0}+1)r}(Y_{i}(X))),$
the right vertical map can be identified with the compositions of following natural maps
[…]
@@
$\longrightarrow K_{*}^{r_{1},\frac{1}{8}}\left(C^{*}_{L,0}\left(P_{s_{0}}\left(\bigsqcup_{j,j^{\prime},X}N_{(s_{0}+1)r_{0}}(Y_{0}^{(j)}(X))\cap N_{(s_{0}+1)r_{0}}(Y_{1}^{(j^{\prime})}(X))\right);\Gamma(Z_{r},\mathcal{A}\otimes\mathcal{K})\right)\right)$
$\longrightarrow K_{*}^{r_{1}s_{1},\frac{1}{8}}\left(C^{*}_{L,0}\left(P_{s_{1}}\left(\bigsqcup_{j,j^{\prime},X}N_{(s_{0}+1)r_{0}}(Y_{0}^{(j)}(X))\cap N_{(s_{0}+1)r_{0}}(Y_{1}^{(j^{\prime})}(X))\right);\Gamma(Z_{r},\mathcal{A}\otimes\mathcal{K})\right)\right)$
$\longrightarrow K_{*}^{(s_{0}+1)r_{1},\frac{1}{8}}\left(I^{\tilde{X},R,\omega,s_{1}}\cap J^{\tilde{X},R,\omega,s_{1}}\right).$
@@
By the coarse invariance and the permanence under taking subspace of the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras, $\partial x$ is sent to 0 for sufficiently large $s_{1}$. By the controlled-exactness of the Mayer-Vietoris sequence, there exists $r_{2}$ and
$y\in K_{*}^{r_{2},\frac{1}{8}}(I^{\tilde{X},R,\omega,s_{1}}),~z\in K_{*}^{r_{2},\frac{1}{8}}(J^{\tilde{X},R,\omega,s_{1}})$
such that $x=y+z$ in $K_{*}^{r_{2},\frac{1}{8}}(B^{\tilde{X},R,\omega,s_{1}})$. Since we have
$I^{X,R,\omega,s_{1}}_{r_{2}}$ $\subset C^{*}_{L,0}\left(P_{s_{1}}\left(\bigsqcup_{j,X}Y_{0}^{(j)}(X)\right)^{+r_{2}};\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K})\right)_{s_{1}r_{2}}$
$\subset C^{*}_{L,0}\left(P_{s_{1}}\left(\bigsqcup_{j}N_{(s_{0}+1)r_{2}}(Y_{0}^{(j)}(X))\right);\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K})\right)_{s_{1}r_{2}}$
$=C^{*}_{L,0}\left(\bigsqcup_{j}P_{s_{1}}\left(N_{(s_{0}+1)r_{2}}(Y_{0}^{(j)}(X))\right);\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K})\right)_{s_{1}r_{2}},$
by the definition (10) of the relative $r_{2}$-neighborhood of $P_{s_{1}}\left(\bigsqcup_{j}Y_{0}^{(j)}(X)\right)$ and by the assumption that each $Y_{0}^{(j)}(X)\in\mathcal{Y}$, $y=0$ in $K_{*}^{r,\frac{1}{8}}(C^{*}_{L,0}(P_{s}(X));\Gamma(\bigsqcup X,\mathcal{A}\otimes\mathcal{K}))$ for sufficiently large $r$ and $s$. Similarly, we have $z=0$ in $K_{*}^{r,\frac{1}{8}}(C^{*}_{L,0}(P_{s}(X));\Gamma(\bigsqcup X,\mathcal{A}\otimes\mathcal{K}))$. Therefore, we have